# Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications

### v3

@@ 3.3 Fine-grained Benchmark Suite
The suite consists of five fine-grained benchmark sets. Each of the five fine-grained benchmarks has been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit compared to the inverse setting. Appendix D contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
@@
Table 3: Accuracy on the Wake Vision (WV) and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
Train
VWW Wake Vision (Quality) WV (Combined)
Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
WV 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
@@
The lighting data sets test the performance of Machine Learning (ML) systems in different lighting conditions. In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine-grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, a simple but effective method for distinguishing lighting conditions [41]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
@@
3.4 Usability
Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [36] and HuggingFace Datasets [22] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
4 Results
In this section, we present a comprehensive evaluation of Wake Vision through five complementary analyses. First, we provide the first insights into best practices for Wake Vision by comparing its two training sets, which also provide initial insights into the trade-off between dataset quality and size in a TinyML setting. Second, we show that Wake Vision serves as an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models. Third, we validate Wake Vision’s benefits across diverse model architectures commonly used in the TinyML community, from MCUNet to ColabNAS families. Fourth, we thoroughly investigate the critical trade-off between dataset quality and size for TinyML models, providing insights into how model capacity influences sensitivity to label noise versus dataset scale. Finally, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios. Throughout these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
4.1 Training Set Evaluation
Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
To evaluate the performance of the two Wake Vision training sets, we train identical models on each set. We use MobileNetV2 [33] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [26]. We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6. After training, we compare each model’s performance on the common test set.
The result of this comparison is shown in Tab. 4. These results show that the Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy. This indicates that label quality is more important than quantity in this setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
4.2 Wake Vision & VWW Cross Evaluation
To validate this drop-in capability and demonstrate Wake Vision’s effectiveness, we do a cross-evaluation study. We train two identical MobileNetv2 [33] models according to the same training recipe as in Sec. 4.1. One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets. Creftab:wv-vww-cross-val shows the result of the cross-evaluation. The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of the Wake Vision (Combined) model described in Sec. 3.1. This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
4.3 Comprehensive Model Architecture Evaluation
Figure 3: Cross-evaluation results on the Wake Vision test set.
Each model is trained on each dataset using the Adam optimizer with a learning rate of 0.001 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set. In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training. We use the checkpoint with the best validation score for testing.
In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision. Results on the Wake Vision test set can be found in Fig. 3, while the results for the VWW test set can be found in Fig. 5 alongside the complete data in Appendix A.
For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW. In some models, the VWW trained variant performs better on the VWW test set. This is only the case for the smallest models in our evaluation suite, which we believe may be due to these models being too small to generalize past the slight domain shift between the VWW and Wake Vision test sets. We additionally see that the accuracy discrepancy between identical models is often highest on Wake Vision’s Test set, further suggesting that the new test set is more challenging.
The boost in accuracy achieved by using the Wake Vision (Quality) Training set can be traded off for smaller model architectures (in some cases 3X fewer MACs) while preserving test accuracy. Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models. Therefore, any boost in accuracy that does not impact the model’s architecture enables further efficiency.
4.4 Impact of Dataset Quality & Size in TinyML
We investigate the impact of dataset quality and size on Resnet [14] style models of varying capacity. We measure quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix E gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [26], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
4.5 Fine-Grained Benchmark Evaluation
Although aggregate metrics like overall accuracy or F1 score provide a high-level view of model performance, they mask critical failure modes and biases that emerge in real-world scenarios. For this reason, we included fine-grained benchmarks in Wake Vision (Section 3.3). Fine-grained evaluation across demographics, environmental conditions, and visual contexts helps us develop reliable TinyML Computer Vision systems for bespoke use cases. This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could have serious implications for fairness and safety.
In Table 5, we compare the models from Sec. 4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
Set Size F1-Score
Benchmark Val Test Wake Vision VWW
Gender Female $684$ $2157$ $0.93$ $0.89$
Male $1310$ $3918$ $0.91$ $0.88$
Unknown $1612$ $4940$ $0.77$ $0.78$
Age Young $275$ $884$ $0.94$ $0.90$
Middle $2133$ $6595$ $0.91$ $0.88$
Older $90$ $276$ $0.94$ $0.89$
Unknown $1299$ $3837$ $0.71$ $0.74$
Distance Near $5457$ 16 3331633316\,33316 333 $0.91$ $0.89$
Medium $2213$ $6876$ $0.85$ $0.84$
Far $398$ $1140$ $0.59$ $0.67$
Lighting Dark $3255$ $9420$ $0.85$ $0.81$
Normal 14 3151431514\,31514 315 43 0104301043\,01043 010 $0.82$ $0.82$
Bright $1012$ $3332$ $0.82$ $0.82$
Depictions Person $356$ $978$ $0.71$ $0.66$
Non-Person $352$ $1101$ $0.86$ $0.82$
No Depiction $8583$ 25 8022580225\,80225 802 $0.87$ $0.85$
The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance dataset at a score of 0.67 and 0.59, respectively. This is likely caused by a difference in how the two datasets are created. Where VWW filters out persons smaller than 0.5% of the image, this filter is set to 5% for Wake Vision. Therefore, the VWW dataset contains images with significantly smaller persons than the Wake Vision dataset, which can improve its performance on far away persons. We did experiment with setting the limit to 0.5% for Wake Vision however, we experienced that it disproportionately decreased the performance of models trained on the dataset in other settings.
5 Ethical Considerations
The Wake Vision dataset and benchmark suite are designed to advance TinyML research by providing data resources and enabling detailed evaluation of person detection systems. While this work has significant potential benefits, we acknowledge several important ethical considerations:
Intended Benefits
Wake Vision can improve the quality of person detection systems that serve beneficial societal purposes, for example, an occupancy detection system that conserves energy by turning off the lights/HVAC/TV when a person leaves. The focus on TinyML and on-device inference helps protect user privacy by avoiding cloud-based processing of sensitive data.
Potential Misuse
We recognize that person detection technology can be misappropriated for harmful purposes, including weapons targeting or mass surveillance systems. While we cannot completely prevent misuse, we implement several safeguards, including releasing only evaluation sets, not training sets, for our fine-grained benchmark datasets.
Data Rights and Privacy
The images in Wake Vision are sourced from Flickr through Open Images [21, 19] under the CC-BY 2.0 license. While we have made extensive efforts to ensure proper licensing, we acknowledge the possibility that some images may have been uploaded without proper distribution rights. Furthermore, the dataset may contain potentially identifying biometric information, and given sufficient resources, individuals could potentially be identified. Due to the scale of the dataset, exhaustive manual verification for offensive content or privacy concerns is infeasible.
Fairness and Representation
Our benchmark suite is specifically designed to evaluate model fairness and robustness across different demographics and conditions. However, we acknowledge that these same demographic benchmarks could be misused to develop systems that classify gender and age without consent, potentially enabling privacy violations or discriminatory practices. We have carefully balanced these concerns against the importance of ensuring that person detection systems work equitably for all populations.
6 Conclusion
The lack of large, high-quality datasets has been a significant bottleneck in TinyML research. Through Wake Vision, we address this challenge by providing a comprehensive dataset and benchmark suite that advances the field in several key dimensions. Wake Vision surpasses the current state-of-the-art VWW dataset by offering  100x more images while maintaining ease of use and improving label accuracy. Our experiments show that Wake Vision serves as an effective drop-in replacement for VWW, consistently improving performance across various model architectures and training scenarios. Moreover, our investigation into dataset quality versus quantity trade-offs reveals important insights specific to TinyML systems. Unlike large models that can effectively learn from noisy data, we find that TinyML models are particularly sensitive to label quality because of their limited capacity. However, we also demonstrate that carefully designed training strategies—such as our approach of pre-training on Wake Vision (Large) followed by fine-tuning on Wake Vision (Quality)—can effectively leverage both quantity and quality to achieve superior performance. Finally, the Wake Vision benchmark suite, with its fine-grained evaluation sets, establishes new standards to evaluate the robustness of the TinyML model in demographics, environmental conditions, and visual contexts. These benchmarks enable researchers to identify and address potential biases and failure modes that could otherwise go undetected using traditional evaluation metrics.
Appendix A Additional Benchmark Results
Table Tab. 6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets. RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”. Links pointing to the original models used for training are also present for each model family involved in the experiment.
Figure Fig. 5 shows cross-evaluation results on the Visual Wake Words test set. The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW. Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
Model input RAM Flash MACs Test Set Training Set
.tflite volume [kiB] [kiB] [MM] WV (Qual.) VWW
MobileNetV2  [^1] https://keras.io/api/applications/mobilenet/[33]
MobileNetV2_0.25 (224,224,3) 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
VWW $88.6\pm 0.17$ $88.3\pm 0.29$
MCUNet  [^2] https://github.com/mit-han-lab/mcunet/blob/master/mcunet/model_zoo.py[24]
10fps_vww (64,64,3) 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
VWW $81\pm 0.12$ $80.8\pm 0.1$
5fps_vww (80,80,3) 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
VWW $83.2\pm 0.57$ $82.6\pm 0.16$
320kb-1mb_vww (144,144,3) 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
VWW $86.5\pm 1.01$ $86\pm 0.25$
Micronets [^3] https://github.com/ARM-software/ML-zoo/tree/master/models/visual_wake_words[2]
vww2_50_50_INT8 (50,50,1) 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
VWW $70.8\pm 0.68$ $70.6\pm 0.64$
vww3_128_128_INT8 (128,128,1) 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
VWW $78.3\pm 0.91$ $77.6\pm 1.13$
vww4_128_128_INT8 (128,128,1) 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
VWW $78.4\pm 0.51$ $76.5\pm 0.37$
ColabNAS444 https://github.com/AndreaMattiaGaravagno/ColabNAS/tree/main/models [11]
k_2_c_3 (50,50,3) 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
VWW $65.6\pm 0.66$ $70.7\pm 0.08$
k_4_c_5 (50,50,3) 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
VWW $69.9\pm 0.26$ $75.5\pm 0,64$
k_8_c_5 (50,50,3) 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
VWW $73\pm 0.91$ $77.3\pm 0.57$
Table 6: Raw data of reconstructed models.
Figure 5: Cross-evaluation results on the Visual Wake Words test set
Appendix B Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
Appendix C Flowchart of Bounding Box Filtering Process
Fig. 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
Appendix D Model Design Case Study
D.1 Image Size vs. Model Width Scaling
D.2 Quantization
D.3 Grayscale
Appendix E Inducing Label Errors
Appendix F Fine Grained Benchmark Images
Fig. 9 gives example images for each of the benchmarks in the suite.

### v5

@@ 3.3 Fine-grained Benchmark Suite
The suite consists of five fine-grained benchmark sets, three of which are applicable to any dataset created by our dataset generation pipeline, while two are specific to Wake Vision. The three generally applicable fine-grained benchmarks include Distance, Lighting, and Depictions, whereas the Wake Vision-specific criteria include Perceived Gender and Perceived Age. Appendix F details all the adopted criteria. Each of the five fine-grained benchmarks has been picked based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting. Appendix E contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
Table 3: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
Train
VWW Wake Vision (Quality) Wake Vision (Combined)
Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
3.4 Community Accessibility and Involvement
Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [45] and HuggingFace Datasets [30] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not shrink due to dead links. The rehosted labels are generated according to our default dataset configuration, further described in Appendix I. Most datasets remain static after the initial release. To continuously improve Wake Vision we partner with the Edge AI Foundation [14] to gather community involvement through recurrent competitions (Sec. 4.5). In our vision, the iterative enhancements brought by the community are integrated into the dataset leading to a continuous improvement cycle.
3.5 Generating Binary Image Classification Datasets for TinyML
Beyond person detection, many other TinyML applications can benefit from large-scale binary classification datasets. For example, bird detection enables smart nest systems [9] and automated feeding  [20], while vehicle detection enables automated parking space occupancy monitoring [6] and blind spot monitoring [43]. We demonstrate that our binary dataset generation pipeline is adaptable for constructing image classification datasets for these and other classes of interest among the 9.6K trainable classes or 600 boxable objects present in Open Images v7 [29, 27].
Constructing binary classification datasets using our pipeline yields significantly larger datasets compared to the prior VWW [7] methodology applied to the same class. For the bird category, the dataset generated via our pipeline contains 12×12$\times$12 × more images than VWW, with a decrease in label error rate from 6.4% to 4.8%. Similarly, for car detection, we provide 27×27$\times$27 × more images with a decrease from 6.6% to 0.6% in label error rate. Further details are available in Appendix S.
4 Results
We present a comprehensive evaluation of Wake Vision through five complementary analyses. First, we compare our two training sets to derive insights on best practices when training. Second, we show that Wake Vision is an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models. Third, we validate Wake Vision’s benefits across diverse model architectures used in the TinyML community, from MCUNet to ColabNAS families. Fourth, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios. Fifth, we present the results of the first Wake Vision Challenge and its resulting community-driven improvements to the Wake Vision dataset. Through these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
4.1 Training Set Evaluation
Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without KD. A MobileNetV2-1.0 model trained on the bounding box set is the teacher.
Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
To evaluate the performance of the two Wake Vision training sets, we train identical models on each set. We use MobileNetV2 [41] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [34], a learning rate of 2e-3 with cosine decay and a weight decay of 4e-6. After training, we compare each model’s performance on the common test set.
The result of this comparison is shown in Table 4. The Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy, indicating that label quality is more important than quantity in this evaluation setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels. We further study how training set size and quality impact accuracy across model sizes, and observe that smaller models are more sensitive to errors in the training set, dropping accuracy by a delta of 0.8% compared to larger models (Appendix C).
4.2 Wake Vision & VWW Cross Evaluation
To validate this drop-in capability, we do a cross-evaluation study. We train two identical MobileNetv2 [41] models according to the same training recipe as in Section 4.1. One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets.
Table 3 shows the result of the cross-evaluation. The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of the Wake Vision (Combined) model described in Section 3.1. This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
4.3 Comprehensive Model Architecture Evaluation
(a) Initial Pareto frontier for Wake Vision vs VWW
(b) Advanced Pareto frontier by challenge submissions
Figure 3: The Wake Vision Challenge submissions advanced our initial Pareto frontiers.
Each model is trained on each dataset using the Adam optimizer with a learning rate of 1e-3 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set. In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training. We use the checkpoint with the best validation score for testing. In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision. Results on the Wake Vision test set can be found in Figure 3(a), while further details and the results for the VWW test set can be found in Appendix A.
For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW. At the extreme, for one Micronets [2] model (middle right of Figure 3(a)), switching the training set to Wake Vision increased the test accuracy by as much as 6.6%. The only instance where Wake Vision isn’t a complete improvement over VWW is in Figure 4 in Appendix A, where small models perform better when trained on the same dataset they are tested on (i.e., trained on VWW and tested on VWW), likely due to ultra-low-capacity models failing to generalize past the slight domain shift between the VWW and Wake Vision test sets.
Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models. Therefore, any boost in accuracy achieved through Wake Vision can also be traded off for more efficient architectures (in some cases 3X fewer MACs) while preserving test accuracy.
4.4 Fine-Grained Benchmark Evaluation
Aggregate metrics like accuracy or F1-score provide high-level insight but can mask critical failure modes and biases that emerge in real-world scenarios. In response, we include fine-grained benchmarks in Wake Vision (Section 3.3) across demographics, environmental conditions, and visual contexts. This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could adversely affect fairness and safety.
In Table 5, we compare the models from Section 4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance benchmark at a score of 0.67 and 0.59, respectively. During dataset generation,VWW filters out persons smaller than 0.5% of the image, versus 5% for Wake Vision. Therefore, VWW contains images with significantly smaller persons. We did experiment with adopting a default of 0.5% for Wake Vision, however this disproportionately decreased the performance of models in other settings.
Table 5: Wake Vision Fine-Grained Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on each benchmark.
Gender Age Distance Lighting Depictions
Female
Male
Unknown
Young
Middle
Older
Unknown
Near
Medium
Far
Dark
Normal
Bright
Person
Non-Person
No Depiction
Size Val 684 1310 1612 275 2133 90 1299 5457 2213 398 3255 14315 1012 356 352 8583
Test 2157 3918 4940 884 6595 276 3837 16333 6876 1140 9420 43010 3332 978 1101 25802
F1 Wake Vision 0.93 0.91 0.77 0.94 0.91 0.94 0.71 0.91 0.85 0.59 0.85 0.82 0.82 0.71 0.86 0.87
VWW 0.89 0.88 0.78 0.90 0.88 0.89 0.74 0.89 0.84 0.67 0.81 0.82 0.82 0.66 0.82 0.85
4.5 The Wake Vision Flywheel: Community-Driven Continuous Improvement
Unlike most datasets which remain static, Wake Vision employs a community-driven flywheel where each competition generates improvements that fuel greater participation and innovation in subsequent rounds, thereby advancing TinyML research and commercial applications. We partner with the Edge AI Foundation [14] to host Wake Vision competitions. The inaugural competition in early 2025 featured two tracks, one focused on developing high-performance, resource-efficient models, and the other on enhancing the Wake Vision person detection dataset. The event garnered significant support from the TinyML community, with 48 participants across both tracks [12].
Successful challenge outcomes included an advanced pareto frontier (Figure 3(b)), containing 3 new models (orange), and a technique to enhance the quality of the Wake Vision large split [4], which in our experiments led to a reduction in the label error rate from 15.2% to 9.8%. Further details can be found in Appendix T. A new competition is currently being launched [13] to further community-led advancements to Wake Vision, and the best results will be integrated in a new version of our dataset. This competition-driven advancement exemplifies our vision of continuous dataset improvement.
5 Ethical Considerations and Limitations
We acknowledge several ethical considerations. Potential Misuse: Our dataset and benchmark suite are designed to advance TinyML research and applications. Intended benefits include occupancy detection systems that conserve energy by turning off the lights/HVAC/TV when a person leaves, and privacy preservation through on-device inference, avoiding cloud processing of camera data. We acknowledge person detection technology can be misappropriated for harmful purposes such as weapons targeting or mass surveillance systems. Data Rights and Privacy: Images are sourced from Flickr through Open Images under the CC-BY 2.0 license. Some images may have been uploaded without proper distribution rights, or may contain potentially identifying biometric information. Fairness and Representation: Our benchmark suite is designed to evaluate model fairness and robustness across different demographics and conditions, ensuring person detection systems work equitably for all populations. However, demographic benchmarks can be misused towards classifying gender and age without consent, potentially enabling privacy violations or discriminatory practices.
6 Conclusion
We introduced an automated binary dataset generation pipeline that addresses a fundamental challenge in TinyML: the systematic creation of large-scale, high-quality datasets tailored to diverse deployment constraints. Our approach combines intelligent multi-source label fusion, confidence-aware filtering, automated label correction, and systematic benchmark generation—innovations necessitated by the prohibitive cost and inconsistency of manual approaches for TinyML’s bespoke dataset requirements. Through Wake Vision, a large-scale binary classification dataset demonstrating our methodology via person detection, we achieved 100$\times$ scale improvement and 6.6% accuracy gains over existing benchmarks. More importantly, we uncovered critical TinyML-specific insights: smaller models exhibit greater sensitivity to label quality than their large-scale counterparts, and optimal performance requires carefully orchestrated two-stage training that balances scale and quality. Our fine-grained evaluation framework revealed performance disparities across demographic and environmental conditions that aggregate metrics fail to capture. The generality of our approach extends beyond person detection, successfully creating production-grade datasets for bird and car detection with similar improvements. By enabling systematic dataset creation across Open Images’ 20K classes, our work provides a methodological foundation that can accelerate TinyML research across diverse application domains. All resources are open-sourced under CC-BY 4.0 license to maximize community impact.
Appendix A Additional Benchmark Results
Table 6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets. RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”. Links pointing to the original models used for training are also present for each model family involved in the experiment.
Figure 4 shows cross-evaluation results on the Visual Wake Words test set. The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW. Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
Table 6: Raw data of reconstructed models.
Model RAM Flash MACs Test Set Training Set
.tflite [kiB] [kiB] [MM] WV (Qual.) VWW
MobileNetV2 [41]
MobileNetV2_0.25 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
VWW $88.6\pm 0.17$ $88.3\pm 0.29$
MCUNet [32]
10fps_vww 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
VWW $81\pm 0.12$ $80.8\pm 0.1$
5fps_vww 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
VWW $83.2\pm 0.57$ $82.6\pm 0.16$
320kb-1mb_vww 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
VWW $86.5\pm 1.01$ $86\pm 0.25$
Micronets [2]
vww2_50_50 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
VWW $70.8\pm 0.68$ $70.6\pm 0.64$
vww3_128_128 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
VWW $78.3\pm 0.91$ $77.6\pm 1.13$
vww4_128_128 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
VWW $78.4\pm 0.51$ $76.5\pm 0.37$
ColabNAS [18]
k_2_c_3 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
VWW $65.6\pm 0.66$ $70.7\pm 0.08$
k_4_c_5 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
VWW $69.9\pm 0.26$ $75.5\pm 0,64$
k_8_c_5 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
VWW $73\pm 0.91$ $77.3\pm 0.57$
Figure 4: Cross-evaluation results on the Visual Wake Words test set
Appendix B Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
Appendix C Impact of Dataset Quality & Size in TinyML
We investigate the impact of dataset quality and size on Resnet [22] style models of varying capacity. We measure quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix C.1 gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [34], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
C.1 Inducing Label Errors
Appendix D Flowchart of Bounding Box Filtering Process
Figure 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
Appendix E Model Design Case Study
E.1 Image Size vs. Model Width Scaling
E.2 Quantization
E.3 Grayscale
Appendix F Fine Grained Benchmark Suite
This section details the fine grained benchmark suite proposed.
@@ Appendix F Fine Grained Benchmark Suite
The lighting data sets test the performance of ML systems in different lighting conditions. In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine-grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, a simple but effective method for distinguishing lighting conditions [50]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
@@
Figure 9 gives example images for each of the benchmarks in the suite.